OpenStudy (anonymous):

\frac{ p+9 }{ 5 }=r for p can somebody help me plz

OpenStudy (anonymous):

Fractions don't show up well in questions... is your question as follows?\[\frac{ p+9 }{ 5 } = r\]

OpenStudy (anonymous):

And you want to solve for "p" ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, great... So, you need to get "p" all alone... first, you have the whole left side of the equation divided by 5. You can't get to the "p" until you get rid of that denominator, so start by multiplying both sides of the equation by 5: (p + 9) / 5 = r (p + 9)*5 / 5 = 5r Then cancel the "5" on the left side: (p+9) = 5r

OpenStudy (anonymous):

okay i got that far im on p+9=r5

OpenStudy (anonymous):

Then, you can get the "p" on the left all alone by subtracting 9 from both sides: (p+9) = 5r p + 9 - 9 = 5r - 9 so... p = 5r - 9

OpenStudy (anonymous):

You probably almost had the whole thing, then :) you were one step away from the answer :)

OpenStudy (anonymous):

yes bit now can u help me with something else

OpenStudy (anonymous):

sure, glad to help... can you post it as a new question?

OpenStudy (anonymous):

yes i can